# ToxicTags: Decoding Toxic Memes with Rich Tag Annotations

### v1

@@ 8 Conclusion
This work makes several key contributions toward advancing research in toxic meme detection and multimodal content moderation. First, we curate a richly annotated dataset of 6,300 real-world meme-based posts through a two-stage labelling process. Second, we enhance contextual understanding by introducing auxiliary metadata including meme titles and, most importantly, social tags. Third, we propose a generative AI-based tag generation module capable of producing socially relevant tags given an arbitrary input meme with no tag information in the first place. Finally, we evaluate the performance of state-of-the-art VLMs under few-shot prompting setups, establishing strong baselines for future research. Collectively, our contributions address key bottlenecks in toxic meme moderation and provide a foundation for building more accurate, context-aware, and socially responsible content moderation systems.
Appendix A Limitations
While this work offers several important contributions, it also has a few limitations. First, our dataset is limited to image-based memes, excluding other emerging modalities such as video and audio, which are increasingly used to disseminate harmful content. Future work will aim to extend this framework to multimodal datasets that better reflect current online communication trends. Second, we do not examine the psychological and emotional impact of hateful memes on viewers. Exposure to such content may contribute to anxiety, depression, or desensitization—an important area that lies beyond the current scope. Addressing this limitation would require collaboration with psychologists, mental health experts, and affected communities. Third, we do not investigate the real-world consequences of online hate memes, particularly their potential to incite offline violence or criminal behavior. Understanding this link between online toxicity and offline harm is a critical direction for future research.
Appendix B Ethics statement
@@ Appendix C Platforms
Data curation platform
As stated earlier, we use imgflip.com as the curation platform for our data since it is centred on conversations using memes. We specifically use the streams [^16] https://imgflip.com/streams feature of the platform to collect these memes. The selected streams and their descriptions are outlined as follows:
  • Dark_Humour ($\sim$11k followers) [^17] https://imgflip.com/m/Dark˙humourWelcome to Dark_humour, Imgflip’s premier community for offensive humour. Stream mood: Relax liberals, it’s called dark humour.
  • Memes_Overload ($\sim$9.5k followers) [^18] https://imgflip.com/m/MEMES˙OVERLOADHey there, welcome to MEMES_OVERLOAD, Imgflip’s 4th largest stream for memes! We’re all here to have fun, so make sure to follow the rules, and keep on memeing on.
@@
Google Lens
To provide contextual information from the real world, we provide two different augmentations to our tag generation module: (i) Google Search API based expansion of tags, and (ii) Google Lens based descriptions. While we have discussed in detail about tag expansion in the main content, here we outline the manual collection procedure and preprocessing applied for Google Lens context incorporation.
Collection
Figure 4: Example of manual collection of title and description of most relevant search obtained from Google Lens.
Figure 4 shows an example meme with corresponding similar results obtained from Google Lens. We manually collect this information for each of the 6,300 memes by strictly asking the collectors to keep only the top-most matching result and to adhere to the following format for information storage: Title – Description. We believe that ours is the first of its kind utilisation of Google Lens for improving contextual knowledge.
Preprocessing
For pre-processing, we follow several steps to clean unwanted textual information from the information returned by Google Lens. First, we remove leading and trailing white spaces. Then we remove certain irrelevant URLs appended within the text, which could have generated ambiguity for the tag generation framework and further went on to clean some platform-specific formats such as subreddit references (e.g. r/, u/). Emojis are eliminated using a Unicode-based regex, and numerical metrics (e.g. 2.5K likes, 3M views) & user mentions (e.g. @user) are also eliminated. Finally, attached metadata that do not contribute to meme semantics and non-English scripts (e.g. Chinese, Arabic, Cyrillic, and Devanagari) are also removed. We use ‘Unicode-range matching’ for targeting only English letters, numbers, and symbols while discarding other languages.
@@ Appendix D Details of annotation
Figure 5 presents a brief pipeline of the employed annotation process over two stages. The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 8. During the whole annotation process, annotators were asked to adhere to the definitions provided in the subsection D.
@@
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
starving cannibal, neon lights,
brothers, cannibalism
Our approach alabama, abortion chocolate, rivers, africa
starving, cannibal,
cannibalism, cannibal
@@
barbecue grill, cook, dog, food,
grill, person, man, meat
RAM++ Openset Close-up Jeans Barbecue chicken, Filling station, Ribs
@@
barbecue, barbecue grill, dog,
food, grill , person, man, meat
RAM Openset ’ ’ Ribs Ribs
@@
anime, girl, people,
@@
mud, puddle, water,
[…]
grill, dog, person, man,
[…]
Table 7: Comparative examples of tags generated by different models.
@@
Tags
signs, advertising
@@
OCR extracted
text
imaflip.com ginger
=== S
@@
BACK RIBS FINGER SANDWHICH
@@
BAKED FRIARS MAC CHEESE
( VEGAN )
@@
GUY : You’re total
is \$ 26.34 ME : I can’t
afford that PIZZA GUY :
Well you’$l$ have to pay
some other way ME :
takes out wallet Wait I
forgot I had 30 dollars
@@
But this is completely
fine HOME TEETH 3
Expert
annotation
hateful dangerous offensive normal
@@
Table 9: Label wise word cloud of tags.
Definitions
hatefulReference(Kiela et al. 2020) – A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. Attack is defined as violent or dehumanizing (comparing people to non-human things, e.g., animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hateful.
@@
offensiveReferences(Roy et al. 2023; Mathew et al. 2022) – A text, meme or speech which is neither hateful, nor dangerous but uses abusive slurs or derogatory terms will be considered as offensive.
toxicReference: PerspectiveAPI [^21] https://developers.perspectiveapi.com/s/about-the-api-model-cards?language=en˙US – A rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.
@@ Appendix E Further analysis of the dataset
sonic . it’s YOUR baby !
@@
MASKS SAVE LINES
@@
tag pairs theme
Table 11: Top co-occurring tag pairs and associated semantic themes.
Related tags: To understand the semantic themes associated with frequently occurring tag pairs, we conduct a co-occurrence analysis. Prior to the analysis, all tags are lemmatised to ensure consistency and improve result accuracy. Our primary objective is to identify how often specific tags co-occur and what thematic contexts they represent within the meme content. The key observations are as follows. Some of the top co-occurring pairs and their themes are listed in Table 11.
@@
Figure 6: Prompts used for Stage I and Stage II classification with few-shot exemplars.
@@ Appendix F Employed models
IDEFICS-3: The latest version of VLM released by HuggingFace team, this is among the few open-source models that are trained with few-shot training objective. We use the instruction fine-tuned version for conducting our experiments – HuggingFaceM4/idefics-9b-instruct.
LLaVA-1.6: An enhanced version of LLaVA (Liu et al. 2023b) and LLaVA-1.5 (Liu et al. 2023a), LLaVA-1.6 has shown significant improvement over its prior models. We use llava-hf/llava-v1.6-mistral-7b-hf checkpoint for running our experiments.
@@
Qwen-2.5 VL: Successor of Qwen VL and Qwen-2 VL, this latest iteration introduces streamlined and efficient vision encoder and dynamic resolution for video understanding. We use the instruction-tuned checkpoint from HuggingFace – Qwen/Qwen2.5-VL-7B-Instruct in our experiments.
GPT-4o: One of the first models by OpenAI team to have multimodal capability, GPT-4o (OpenAI 2024) has proved its wide applicability across multiple domains [^23] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
@@
Appendix G Employed prompts
Detailed list of prompts for the corresponding variants for Stage-I and Stage-II are provided 6.
Appendix H Experimental setup
(i) Experimental setup for tag generation module:
(a) For generating caption using the LaMa inpainted image, we set max new tokens to 30.
(b) For generating tagged summary we have used maximum number of tokens to 150.
(c) For finetuned summary generator and tag generation module, all the models are finetuned using Trainer API, where we set batch size as 1, gradient accumulation steps set to 1 and a learning rate of 2e-5. We use AdamW Torch optimizer with a $\beta_{2}$ value of 0.999 and a weight decay of 1e-6. Transformer modules including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj and down_proj, with a rank value of 8 and task type set to causal language modeling, are finetuned using LoRA-based parameter-efficient fine-tuning. We use 4-bit quantization with bfloat16 compute precision through the BitsAndBytesConfig for efficient memory usage and faster training. To stabilize the initial learning phase, we set a warmup step of two and the number of training epochs to three. To control the maximum sequence length, we set max length to 512, while truncation and padding are also applied.
(ii) Few-shot prompting: To control the randomness of the generated output, we set the temperature to 0.001 for all models used in Stage I and II. The max new tokens was set to 30 and all other parameters were kept as default.

### v2

@@ 9 Conclusion
This work makes several key contributions toward advancing research in toxic meme detection and multimodal content moderation. First, we curate a richly annotated dataset of 6,300 real-world meme-based posts through a two-stage labelling process which we call ToxicTags. Second, we enhance contextual understanding by introducing auxiliary metadata including meme titles and, most importantly, social tags. Third, we propose StemTox that detects toxicity as well as generates tags given an arbitrary input meme. Finally, we compare the performance of toxicity detection and tag generation on an array of benchmarks and baselines for generalizability. Collectively, our contributions address key bottlenecks in semantically aligned toxic meme moderation and thus provides a foundation for building more accurate, context-aware, and socially responsible systems.
10 Limitations
While this work offers several important contributions, it also has a few limitations. First, our dataset is limited to image-based memes, excluding other emerging modalities such as video and audio, which are increasingly used to disseminate harmful content. Future work will aim to extend this framework to multimodal datasets that better reflect current online communication trends. Second, we do not examine the psychological and emotional impact of hateful memes on viewers. Exposure to such content may contribute to anxiety, depression, or desensitization – an important area that lies beyond the current scope. Addressing this limitation would require collaboration with psychologists, mental health experts, and affected communities. Third, we do not investigate the real-world consequences of online hate memes, particularly their potential to incite offline violence or criminal behaviour. Understanding this link between online toxicity and offline harm is a critical direction for future research.
Appendix A Ethics statement
@@
Appendix B Prompt templates
We display below the prompts used for Stage I and Stage II classifications. To generate the image caption, we first apply LLaMA image inpainting  Suvorov et al. (2021) to remove any OCR text from the image, and then prompt GPT to produce a caption.
System prompt for stage I You are an AI assistant tasked with classifying memes.
Consider the following definitions.
$\bullet$ toxic: {definition of toxic}
$\bullet$ normal: {definition of normal}
Based on the above definitions, the input image, title, provided caption, list of tags and the extracted OCR text from the image delimited by three backticks classify the meme as toxic or normal.
Example output for toxic meme: {toxic.}
Example output for normal meme: {normal}
System prompt for stage II You are an AI assistant tasked with classifying memes.
Consider the following definitions.
$\bullet$ hateful: {definition of hatful}
$\bullet$ dangerous: {definition of dangerous}
$\bullet$ offensive: {definition of offensive}
$\bullet$ normal: {definition of normal}
Based on the above definitions, the input image, title, provided caption, list of tags and the extracted OCR text from the image delimited by three backticks classify the meme as hateful, dangerous, offensive or normal.
Example output for hatful meme: {hateful.}
Example output for dangerous meme: {dangerous.}
Example output for offensive meme: {offensive.}
Example output for normal meme: {normal}
@@ Appendix C Platforms
C.1 Data curation platform
As stated earlier, we use https://imgflip.com as the curation platform for our data since it is centred on conversations using memes. We specifically use the streams [^14] https://imgflip.com/streams feature of the platform to collect these memes. The selected streams and their descriptions are outlined as follows:
  • Dark_Humour ($\sim$11k followers) [^15] https://imgflip.com/m/Dark_humourWelcome to Dark_humour, Imgflip’s premier community for offensive humour. Stream mood: Relax liberals, it’s called dark humour.
  • Memes_Overload ($\sim$9.5k followers) [^16] https://imgflip.com/m/MEMES_OVERLOADHey there, welcome to MEMES_OVERLOAD, Imgflip’s 4th largest stream for memes! We’re all here to have fun, so make sure to follow the rules, and keep on memeing on.
@@ Appendix D Details of annotation
Figure 1 presents a brief pipeline of the employed annotation process over two stages. The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 10. During the whole annotation process, annotators were asked to adhere to the definitions provided in the subsection D.1.
@@
Image [Uncaptioned image] [Uncaptioned image]
@@
StemTox alabama, problems, abortion, pregnancy chocolate, rivers, africa
@@
RAM++ (O) Close-up Jeans
@@
RAM (P)
RAM (O) Ribs
@@
anime, girl, people, person
User Specified
mud, puddle, water, man
User Specified
@@
Table 9: Comparative examples of tags generated by different models. (O): open-set, (P): pre-trained using ToxicTags.
@@
Tags signs, advertising
@@
OCR extracted text imaflip.com ginger === S
@@
BACK RIBS FINGER SANDWICH
@@
BAKED FRIARS MAC CHEESE (VEGAN)
@@
GUY: Your total is \$26.34
ME: I can’t afford that
PIZZA GUY: Pay another way
@@
But this is completely fine
HOME TEETH 3
Expert annotation hateful dangerous offensive normal
@@
Table 11: Label-wise word cloud of tags.
D.1 Definitions
hatefulReferenceKiela et al. (2020) – A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. Attack is defined as violent or dehumanizing (comparing people to non-human things, e.g., animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hateful.
@@ D.1 Definitions
offensiveReferencesRoy et al. (2023); Mathew et al. (2022) – A text, meme or speech which is neither hateful, nor dangerous but uses abusive slurs or derogatory terms will be considered as offensive.
toxicReference: PerspectiveAPI [^19] https://developers.perspectiveapi.com/s/about-the-api-model-cards?language=en_US – A rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.
@@ Appendix E Further analysis of the dataset
sonic, it’s YOUR baby!
@@
MASKS SAVE LIVES
Ground truth
annotation
dangerous dangerous dangerous
@@ Appendix F Employed models
LLaVA-1.5: LLaVA-1.5 Liu et al. (2023) has shown significant improvement over its prior models. We use llava-hf/llava-1.5-7b-hf checkpoint for running our experiments.
GPT-4o: One of the first models by OpenAI team to have multimodal capability, GPT-4o OpenAI (2024) has proved its wide applicability across multiple domains [^21] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
@@
Appendix G Experimental setup
We train both, uni- and multi-tasking framework using a consistent experimental configuration to ensure a fair comparison. For multi-tasking, the training objective consists of two components: a weighted cross-entropy loss for the classification task and the default autoregressive generation loss for toxic tag generation. For uni-tasking there is no generation loss. For computing class weights, we follow the weighting scheme described in King and Zeng (2001). All input sequences are truncated or padded to a fixed maximum sequence length of 2048 tokens, and the random seed is fixed to 42 to ensure reproducibility.
Training is performed for three epochs with a per-device batch size of 1 and a gradient accumulation step of 4, resulting in an effective batch size of 4. We use the AdamW optimizer with a learning rate of $2\times 10^{-5}$, weight decay of $1\times 10^{-6}$, and $\beta_{2}$ set to 0.999, along with a warm-up of two steps. For memory-efficient training, the model is loaded using 8-bit quantization via the BitsAndBytes configuration, with computations carried out in FP16 precision. For the multi-task Learning framework, we apply LoRA-based parameter-efficient fine-tuning with rank 8, targeting the following modules: {q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj}.
In contrast, for the uni-task learning framework, the model backbone is fully frozen during training. Extracted hidden representations are normalized using the default normalization layers provided by the backbone model at the time of forward propagation. A sequence compression module is used to extract features into a fixed-dimensional embedding, which is then passed through a label classification head to predict the final class labels.
For the ToxicTags dataset, we fix the number of classification classes to four and for the FHM and MAMI datasets, we fix the number of classes to two. For result calculation and analysis of stage I, we merge the hateful, dangerous, and offensive categories into a single toxic class. This aggregation is the reverse of the two-stage annotation strategy followed during dataset construction.